# Toward a unified theory for common affine roots of general sets of multivariate polynomials

### v2

@@ 5 Concluding remarks
In this paper we proposed a unified theory for common affine roots of general sets of multivariate polynomials over any field, except we did not treat multiplicity. For multivariate polynomials there are several concepts of multiplicity but in particular the one related to Hasse-derivatives seems to have gained interest in the literature [38, 34, 13, 39, 15, 26, 14]. We leave it for further research to investigate if the results in there can be further developed to possibly obtain sharper findings along the line of Theorem 5, Corollary 10 and Remark 11.
We further pose it as a research problem to enhance our findings to hold for integral domains, skew fields and possibly non-commutative integer rings generalizing results in [7, 35]. Also we believe that the material of Section 3 should have implications for the study of particular types of point-sets such as roots of absolutely irreducible polynomials etc. Similarly it would be interesting to see if some of the methods in the present paper can be applied to treat projective points beyond what is already in the litterature e.g. in [11, 5, 33]
We note that like many other recent results on roots of sets of multivariate polynomials also the present paper is inspired by work in coding theory and cryptography. For this particular paper the inspiration comes from work on ramp secret sharing schemes [22].

### v3

@@ 5 Concluding remarks
In this paper we proposed a unified theory for common affine roots of general sets of multivariate polynomials over any field, except we did not treat multiplicity. For multivariate polynomials there are several concepts of multiplicity and for more of those a footprint bound has been formulated [11, Ch. 4, Cor. 2.5] and [39, 26] as well as other types of bounds [16, 29]. To the best of our knowledge not much has been done regarding the interpolation problem (II) taking multiplicity into account. We leave it for further research to establish such results along the line of Theorem 5 and to link them to the above mentioned solutions of (I) in a fashion similar to Remark 11.
We further pose it as a research problem to enhance our findings to hold for integral domains following up on [7, 40] and possibly also to treat skew fields and other non-commutative algebraic structures where the OWB point of view in Section 3 may prove essential. Similarly it would be interesting to see if some of the methods in the present paper could be applied to treat projective points beyond what is already in the litterature e.g. in [12, 5, 37].
We note that like other recent results on roots of sets of multivariate polynomials, e.g. [17], also the present paper is related to work in coding theory and cryptography. For this particular paper the inspiration came from work on ramp secret sharing schemes [24]. Elaborating further on this relation we observe, that Theorem 17 should have interesting implications for one-point algebraic geometric codes [43, Sec. II] and other code constructions where one can obtain information on the functions $\bar{\sigma}$ as well as $\bar{\mu}$.